OpenStudy (anonymous):

Could someone help me with this? Find the domain of the given function, and express the domain in interval notation. g(t) = t^2 = 3t

OpenStudy (anonymous):

edit: g(t) = t^2 + 3t I wasn't sure if I should set it to t^2 + 3t > or = 0?

OpenStudy (anonymous):

you have to many equal signs, but the domain of any polynomial is all real numbers you can write \(\mathbb{R}\) or \((-\infty,\infty)\)